a man's age is triple the age of his son. in eleven years the man's age would be the double of his son's age then. what are the present ages of the man and the son?
Answer by TimothyLamb(4379)

x = man age now
y = son age now
---
x = 3y
x + 11 = 2(y + 11)
x + 11 = 2y + 22
---
put the system of linear equations into standard form
---
x - 3y = 0
x - 2y = 11
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 33
y = 11
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---